OpenStudy (anonymous):

the major axis of an ellipse is horizontal and its length is 24 the length of the minor axis is 14 and the center is (3,-8) find the equation of the ellipse in standard form

OpenStudy (anonymous):

So a = 12 because the major axis =2a and b=7 because the minor axis = 2b. Since this is a horizontal ellipse, it will be in the form (x-h)^2 / a^2 + (y-k)^2 /b^2 =1. So h=3, and k=-8. (x-3)^2 /144 + (y+8)^2 /49 =1 :) Hope this helps!